# RS-RAG: Bridging Remote Sensing Imagery and Comprehensive Knowledge with a Multi-Modal Dataset and Retrieval-Augmented Generation Model

### v1

@@ V-C Results on Image Classification Task
Table IV reports the performance of all models on the image classification task using a subset of the RSWK dataset, measured by overall accuracy and per-class accuracy across 15 scene categories. Among the baselines, LLaMA-3.2-Vision-11B achieves the highest overall accuracy at 34.0 percent, followed closely by InternVL2.5-8B and Qwen2.5-VL-7B. Despite these results, baseline models perform poorly on several knowledge-dependent categories such as Church, Mansion, Historic Site, and Museum, indicating their limited capacity to reason about semantically nuanced or infrequent classes. Our RS-RAG-11B model achieves a substantial improvement, attaining 84.2 percent overall accuracy. It outperforms all baselines by a large margin across nearly all categories, including challenging ones like Church with 82.0 percent accuracy and Historic Site with 65.5 percent accuracy. RS-RAG-7B also surpasses all baselines, reaching 65.9 percent overall accuracy, despite using a smaller backbone. These results demonstrate the effectiveness of integrating retrieved world and domain knowledge, which helps the model resolve semantic ambiguities and enhances recognition of visually similar or context-dependent categories in remote sensing imagery. To better understand the effectiveness of our RS-RAG framework, we conduct a qualitative comparison of image classification outputs generated by baseline models and our RS-RAG variants, as shown in Fig. 5. The figure presents two representative examples from the RSWK dataset: the Denver Museum of Nature and Science and the Bahrain World Trade Center. In both cases, existing vision-language baselines misclassify the scene, assigning incorrect labels such as ”University,” ”Government Building,” or ”Historic Site.” In contrast, RS-RAG-7B and RS-RAG-11B correctly identify the ground-truth categories, demonstrating a better understanding of scene semantics. These results qualitatively highlight the strength of retrieval-augmented generation in capturing contextual cues and leveraging external knowledge to improve fine-grained classification in complex remote sensing scenarios.
@@
TABLE V: Performance comparison between baseline models and our proposed RS-RAG variants on the image VQA task using a subset of the RSWK dataset
@@
InternVL2.5-8B [29] 0.170 0.106 0.073 0.051 0.130 0.178 0.204
Janus-Pro-7B [30] 0.101 0.060 0.039 0.025 0.113 0.124 0.084
Qwen2.5-VL-7B [31] 0.116 0.067 0.042 0.026 0.124 0.135 0.105
Llama3.2-Vision-11B [32] 0.092 0.056 0.037 0.024 0.114 0.115 0.064
RS-RAG-7B 0.233 0.179 0.144 0.113 0.214 0.258 0.608
RS-RAG-11B 0.250 0.197 0.161 0.128 0.249 0.266 0.536
@@ V-D Results on Visual Question Answering Task
Table V presents the quantitative evaluation of various models on the VQA task using a subset of the RSWK dataset. Across all metrics—including BLEU, METEOR, ROUGE-L, and CIDEr—the proposed RS-RAG variants outperform all baselines by substantial margins. RS-RAG-11B achieves the highest scores on BLEU-1, BLEU-4, and METEOR, indicating more fluent and semantically relevant answers. RS-RAG-7B also delivers competitive results, achieving the best CIDEr score of 0.608, suggesting strong alignment with human-annotated answers in terms of informativeness and precision. In contrast, baseline models such as Qwen2.5-VL-7B and Janus-Pro-7B struggle to generate accurate or coherent responses, especially in knowledge-intensive queries. To better understand the sources of these improvements, Fig. 6 illustrates two representative qualitative examples. In the first case, the model is asked about the seating capacity of a stadium. Baseline responses are either factually incorrect or overly generic, whereas both RS-RAG models retrieve and accurately generate the correct capacity of 30,118. In the second example, involving a historically grounded question, only RS-RAG models correctly identify Michael Jackson’s concerts at Burswood Dome during his HIStory World Tour in 1996. Other models either hallucinate unrelated content, such as the 1996 Summer Olympics, or fail to recognize the event entirely. These examples underscore the strength of RS-RAG in incorporating relevant domain and world knowledge to produce accurate, informative, and context-aware answers in remote sensing VQA tasks.
V-E Ablation Studies
Effect of the number of retrieved candidates. To investigate the impact of the number of retrieved candidates in the retrieval-augmented generation process, we conduct an ablation study on the top-$k$ parameter in RS-RAG, as shown in Table VI. With the fusion weight $\alpha$ in Eq.7 fixed to 0.9, we evaluate the model’s performance on the image captioning task under different values of $k$. The results show that retrieving a single, highly relevant knowledge snippet ($k=1$) yields the best performance across all evaluation metrics. As $k$ increases to 3 and 5, the performance consistently degrades, indicating that incorporating more candidates introduces semantic redundancy or irrelevant noise. Such noise can interfere with the model’s ability to generate concise and accurate captions. These findings underscore the importance of retrieval precision in the remote sensing domain and confirm that fewer, but higher-quality, knowledge segments are more effective for guiding the generation process.
@@ V-E Ablation Studies
TABLE VI: Ablation study on the number of top-$k$ retrieved candidates for RS-RAG.
@@
$k=1$ 0.562 0.449 0.385 0.336 0.308 0.362 0.072
$k=3$ 0.458 0.339 0.278 0.232 0.257 0.283 0.044
$k=5$ 0.433 0.314 0.255 0.210 0.244 0.271 0.063
Effect of fusion weight $\alpha$. We further explore the effect of the fusion weight $\alpha$ that balances visual and textual similarity in the re-ranking step (see Eq.7). TableVII reports performance under different values of $\alpha$, with the number of retrieved candidates fixed to $k=1$. The results show that $\alpha=0.5$ achieves the best overall performance, indicating that equal weighting of visual and text similarity provides the most informative guidance for downstream caption generation. When $\alpha$ is too low (e.g., 0.3), text similarity dominates, leading to less visually grounded results. Conversely, when $\alpha$ increases beyond 0.5, performance slightly degrades, suggesting that over-reliance on visual similarity may overlook relevant textual semantics. These findings confirm that effective cross-modal fusion is critical for high-quality retrieval and generation.
TABLE VII: Ablation study on the fusion weight $\alpha$ in RS-RAG.
@@
0.3 0.414 0.253 0.179 0.127 0.215 0.199 0.005
0.5 0.582 0.473 0.410 0.361 0.318 0.393 0.210
0.7 0.568 0.454 0.389 0.338 0.309 0.366 0.110
0.9 0.562 0.449 0.385 0.336 0.308 0.362 0.072
@@ VI Conclusion
In this work, we presented RS-RAG, a retrieval-augmented vision-language framework designed to bridge remote sensing imagery with structured domain and world knowledge. To support this framework, we constructed the RSWK dataset, a large-scale multimodal benchmark that integrates high-resolution satellite imagery with rich textual descriptions covering over 14,000 globally recognized locations from 175 countries. By leveraging this curated knowledge base, RS-RAG significantly improves contextual reasoning and semantic understanding across key vision-language tasks, including image captioning, image classification, and visual question answering. Extensive experiments validate the effectiveness of our approach, particularly in handling complex, knowledge-intensive queries. We believe that both the RSWK dataset and the RS-RAG framework provide a strong foundation for advancing research in remote sensing vision-language understanding and knowledge-grounded geospatial AI.

### v2

@@ V-C Results on Image Classification Task
We further evaluate the capabilities of vision-language models on the image classification task using the RSWK-Mini dataset. As shown in Table V, RS-RAG significantly outperforms all baseline models in terms of overall accuracy as well as performance across most individual categories. Specifically, RS-RAG achieves an overall accuracy of 0.79, representing a substantial improvement over the strongest baseline, Qwen2.5-VL, which achieves 0.61. In terms of category-specific performance, RS-RAG consistently delivers top results in 13 out of 16 categories, with particularly high accuracies for Airport (0.98), Stadium (0.99), Mansion (0.92), and Park (0.80). For commonly encountered scenes in remote sensing, such as Forest, although RS-RAG does not achieve the highest score, its accuracy of 0.89 still demonstrates strong capability in handling typical geospatial scenarios. These results underscore the effectiveness of our retrieval-augmented design in enhancing the scene recognition capacity of VLMs, particularly in remote sensing contexts that demand fine-grained semantic understanding. To further evaluate RS-RAG’s capability in scene classification, we conduct a qualitative comparison on two challenging urban examples, as shown in Fig. 5. The task requires selecting the correct category from 16 visually and semantically similar classes. In the first case, RS-RAG correctly classifies the Hamburg Rathaus as a Government Building, while all baseline models fail, offering incorrect labels like “Museum” or “Amusement Park.” In the second case, only RS-RAG accurately identifies the Brukenthal National Museum, whereas other models misclassify it as a Church or Historic Site. These results demonstrate RS-RAG’s superior ability to align visual content with structured knowledge, enabling robust performance in semantically ambiguous remote sensing scenes.
@@
TABLE VI: Performance comparison between baseline models and our proposed RS-RAG variants on the image VQA task using the RSWK-Mini dataset
@@
InternVL2.5 [6] 0.095 0.038 0.024 0.018 0.078 0.182 0.134
Janus-Pro [5] 0.094 0.038 0.024 0.017 0.076 0.180 0.131
Qwen2.5-VL [31] 0.102 0.041 0.027 0.019 0.083 0.187 0.158
Llama3.2-Vision [1] 0.102 0.042 0.027 0.019 0.082 0.187 0.153
RSGPT [10] 0.209 0.132 0.097 0.072 0.201 0.347 0.569
Geochat [13] 0.207 0.129 0.093 0.068 0.193 0.353 0.531
RS-RAG 0.254 0.149 0.108 0.080 0.239 0.375 0.685
@@ V-D Results on Visual Question Answering Task
To assess the models’ ability to handle knowledge-intensive multimodal reasoning, we conduct evaluations on the VQA task using the RSWK-Mini dataset. As shown in Table VI, RS-RAG delivers superior performance across all evaluation metrics. It achieves a BLEU-1 score of 0.254 and a CIDEr score of 0.685, significantly surpassing the strongest baseline, RSGPT, by 4.5 and 11.6 points, respectively. Although RSGPT exhibits relatively strong performance, RS-RAG consistently outperforms it across all metrics, especially in higher-order BLEU scores and METEOR, which reflect deeper semantic understanding and language fluency. These results underscore the effectiveness of our retrieval-augmented strategy in integrating both remote sensing knowledge and structured world knowledge into the vision-language reasoning process. The consistent improvements across BLEU, METEOR, ROUGE-L, and CIDEr indicate that RS-RAG not only retrieves relevant content but also generates contextually coherent and semantically meaningful answers.
To further illustrate the model’s effectiveness, we provide qualitative examples of the three types of VQA tasks in Fig. 6. RS-RAG consistently generates answers that are not only factually accurate but also grounded in structured remote sensing and world knowledge. In the first example, which corresponds to comprehensive reasoning, RS-RAG correctly interprets the impact of low albedo on the urban heat island effect by incorporating retrieved scientific knowledge, whereas other models either hallucinate surface reflectance values or fail to link them to urban climate phenomena. In the second example, which tests remote sensing knowledge, RS-RAG accurately interprets the meaning of a low NDBI value as an indicator of low urbanization, whereas most baselines confuse land cover classification with index interpretation. In the third example, which focuses on world knowledge, RS-RAG identifies both the historical event and its relevance to the landmark, while other models misattribute unrelated events or fail to provide temporally aligned responses. These results highlight the strength of RS-RAG in connecting visual content with precise, knowledge-grounded answers across different question types, demonstrating its robustness in addressing complex multimodal reasoning within the remote sensing domain.
V-E Discussion
1) Effect of the number of retrieved candidates. To understand the influence of retrieval granularity on downstream performance, we conduct an ablation study by varying the number of top-$k$ retrieved candidates in RS-RAG across three tasks: image captioning, image classification, and VQA, as presented in Table VII. As shown in Table VII (a) for image captioning, setting $k$ to 1 already yields the best performance across most evaluation metrics. Specifically, RS-RAG achieves a BLEU-1 score of 0.490, a METEOR score of 0.447, and a ROUGE-L score of 0.476. When $k$ increases to 3 and 5, performance in BLEU and METEOR declines slightly, while CIDEr improves marginally, suggesting that incorporating additional retrieved knowledge may introduce noise while marginally enhancing informativeness. For image classification, shown in Table VII (b), a consistent degradation in performance is observed as $k$ increases. The best overall accuracy and CIDEr score are obtained when $k$ is 1, indicating that precise and highly relevant knowledge is essential for accurate scene categorization, whereas retrieving more candidates may introduce irrelevant or distracting context. In contrast, VQA results presented in Table VII (c) show a clear benefit from broader retrieval. Performance steadily improves as $k$ increases, with the highest scores observed when $k$ is 5. Under this setting, RS-RAG achieves a BLEU-1 score of 0.254, a METEOR score of 0.239, and a CIDEr score of 0.685. This suggests that answering complex and knowledge-intensive questions benefits from access to a wider and more diverse set of evidence. Overall, these results highlight the importance of task-specific retrieval strategies. While minimal retrieval is sufficient for captioning and classification, broader retrieval is critical for effective multimodal reasoning in VQA. The consistent performance of RS-RAG across varying retrieval settings underscores its robustness and adaptability in handling diverse remote sensing tasks with structured knowledge.
@@ V-E Discussion
TABLE VII: Ablation study on the number of top-$k$ retrieved candidates for RS-RAG across three tasks: image captioning, classification, and VQA.
(a) Image Captioning
@@
$k=1$ 0.490 0.366 0.300 0.252 0.447 0.476 0.145
$k=3$ 0.483 0.358 0.293 0.245 0.439 0.472 0.147
$k=5$ 0.485 0.362 0.297 0.250 0.441 0.474 0.173
(b) Image Classification
Top-$k$ BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
$k=1$ 0.794 0.358 0.213 0.163 0.466 0.795 2.376
$k=3$ 0.676 0.289 0.175 0.136 0.387 0.676 1.966
$k=5$ 0.654 0.272 0.167 0.130 0.369 0.656 1.875
(c) Image VQA
Top-$k$ BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
$k=1$ 0.242 0.136 0.097 0.071 0.220 0.359 0.594
$k=3$ 0.246 0.140 0.100 0.074 0.228 0.365 0.617
$k=5$ 0.254 0.149 0.108 0.080 0.239 0.375 0.685
2) Effect of fusion weight $\alpha$. We further investigate the impact of the modality fusion weight $\alpha$ on downstream performance by varying its value across three tasks: image captioning, image classification, and VQA. As summarized in Table VIII, this coefficient controls the balance between visual and textual similarity scores in the re-ranking stage of retrieval. In the image captioning task, performance improves consistently as $\alpha$ increases. The highest scores are obtained when $\alpha$ is set to 0.9, yielding a BLEU-1 score of 0.490, a METEOR score of 0.447, and a CIDEr score of 0.145. These results suggest that visual similarity plays a dominant role in generating accurate and contextually appropriate captions for remote sensing imagery. Interestingly, the image classification task exhibits a different pattern. The best performance is achieved at $\alpha$ set to 0.5, where RS-RAG obtains a BLEU-1 score of 0.748, a METEOR score of 0.433, and a CIDEr score of 2.376. This indicates that a more balanced integration of visual and textual modalities is critical for effective scene categorization, possibly due to the complementary nature of semantic and spatial cues. In contrast, the VQA task again favors higher $\alpha$ values. When $\alpha$ is set to 0.9, the model achieves its best performance, with BLEU-1 of 0.246, METEOR of 0.228, and CIDEr of 0.617. These results highlight the importance of visual evidence in answering spatially grounded or visually intensive questions. Overall, these findings underscore the necessity of task-specific fusion strategies. While image captioning and VQA benefit from emphasizing visual similarity, image classification requires a more nuanced balance between modalities. This flexibility further demonstrates the robustness and adaptability of the RS-RAG framework in knowledge-driven remote sensing scenarios.
TABLE VIII: Ablation study on the fusion weight $\alpha$ in RS-RAG across three tasks: image captioning, classification, and VQA.
(a) Image Captioning
@@
0.3 0.364 0.218 0.153 0.113 0.311 0.317 0.024
0.5 0.399 0.256 0.190 0.147 0.347 0.355 0.066
0.7 0.477 0.347 0.280 0.231 0.430 0.453 0.166
0.9 0.490 0.366 0.300 0.252 0.447 0.476 0.145
(b) Image Classification
$\alpha$ BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
0.3 0.669 0.302 0.179 0.137 0.393 0.669 2.004
0.5 0.748 0.329 0.197 0.152 0.433 0.748 2.208
0.7 0.711 0.309 0.186 0.144 0.410 0.711 2.087
0.9 0.676 0.289 0.175 0.136 0.387 0.676 1.966
(c) Image VQA
$\alpha$ BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
0.3 0.213 0.122 0.088 0.064 0.209 0.342 0.507
0.5 0.212 0.121 0.087 0.064 0.208 0.339 0.506
0.7 0.211 0.121 0.087 0.064 0.208 0.338 0.503
0.9 0.246 0.140 0.100 0.074 0.228 0.365 0.617
TABLE IX: Performance comparison between baseline models and our proposed RS-RAG variants on the image captioning task using the ChatEarthNet dataset.
Model BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
InternVL2.5 [6] 0.369 0.213 0.134 0.082 0.302 0.300 0.076
Janus-Pro [5] 0.366 0.210 0.131 0.080 0.306 0.295 0.070
Qwen2.5-VL [31] 0.370 0.209 0.126 0.074 0.308 0.284 0.079
Llama3.2-Vision [1] 0.370 0.214 0.136 0.085 0.306 0.304 0.077
RSGPT [10] 0.370 0.217 0.136 0.084 0.295 0.299 0.087
Geochat [13] 0.335 0.201 0.128 0.078 0.301 0.284 0.016
RS-RAG 0.388 0.243 0.153 0.098 0.351 0.356 0.171
3) Extension of RS-RAG on ChatEarthNet Dataset. To further validate the generalization capability of RS-RAG beyond the RSWK dataset, we conduct additional experiments on the ChatEarthNet dataset [34], a recently proposed large-scale image–text dataset designed for training large vision–language models. As shown in Table IX, RS-RAG consistently outperforms all baseline models across all seven evaluation metrics, including BLEU-1 to BLEU-4, METEOR, ROUGE-L, and CIDEr. Specifically, RS-RAG achieves a BLEU-1 score of 0.388 and a CIDEr score of 0.171, marking respective gains of 1.8 and 8.4 points over the strongest baseline, RSGPT. These improvements underscore the robustness of RS-RAG in adapting to new remote sensing datasets and generating semantically rich, knowledge-grounded captions. The consistent performance advantage across metrics further confirms the effectiveness of our retrieval-augmented design in enhancing vision–language understanding, even in domains beyond the original dataset.
4) Generalization in Visually Similar Scenes. To further evaluate the generalization capability of RS-RAG beyond landmark recognition, we conducted additional experiments on common scenes that exhibit high visual similarity. Specifically, we collected 1,000 samples each for grassland and farmland from diverse global regions and compared RS-RAG with the state-of-the-art VLM Qwen2.5-VL. These two categories share highly similar visual patterns in satellite imagery, making them challenging to distinguish based solely on visual cues. As shown in Table X, RS-RAG consistently outperforms the baseline across both categories, demonstrating its superior reasoning capability when visual information is insufficient.
TABLE X: Comparison of classification accuracy on visually similar categories: Grassland and Farmland.
Category Qwen2.5-VL RS-RAG Improvement
Grassland 0.834 0.904 +7.0%
Farmland 0.955 0.990 +3.5%
In addition to the quantitative comparison on grassland and farmland, we further collected 1,000 urban-block samples and prompted the models to identify their corresponding geographic regions. As shown in Fig. 7, RS-RAG effectively recognizes subtle urban contexts by integrating retrieved geographic and environmental knowledge, whereas the Qwen2.5-VL often confuses visually similar structures. For instance, the baseline misclassifies Xicheng District, Beijing as Old Delhi, India due to architectural resemblance, while RS-RAG correctly infers the location using contextual cues such as coordinates and population density. A similar improvement is observed when distinguishing Lehigh Valley region, Pennsylvania from Baltimore metropolitan area, where RS-RAG exhibits more precise spatial reasoning grounded in external knowledge. These results demonstrate that RS-RAG offers clear advantages in fine-grained, visually ambiguous, and common-scene reasoning tasks, underscoring the necessity of retrieval augmentation beyond traditional visual recognition.
@@ VI Conclusion
In this work, we presented RS-RAG, a retrieval-augmented vision-language framework designed to bridge remote sensing imagery with structured domain and world knowledge, thereby addressing the challenges of knowledge-intensive multimodal reasoning. To support this framework, we constructed the RSWK dataset, a large-scale, high-resolution multimodal benchmark that includes imagery-text pairs from over 14,000 geographically and semantically diverse landmarks, spanning 16 categories across 184 countries. Building upon RSWK, we established a comprehensive benchmark covering three representative tasks, including image captioning, image classification, and VQA. To better exploit the dataset’s rich semantics, we further designed the VQA task into three distinct subtypes, targeting comprehensive reasoning, remote sensing knowledge understanding, and world knowledge understanding. Extensive experiments conducted on the RSWK benchmark and the external multimodal dataset demonstrate that RS-RAG consistently outperforms strong baselines across multiple tasks, with notable advantages in handling knowledge-intensive queries that require both factual accuracy and sequential reasoning. We believe that both the RSWK dataset and the RS-RAG framework establish a new foundation for advancing knowledge-grounded vision-language understanding in remote sensing and geospatial AI. Our work paves the way toward more capable, explainable, and generalizable multimodal systems for Earth observation and related applications.